OpenStudy (anonymous):

find an equation of the line passing through the point (5,4) and parallel to the line whose equation is 2x+y=3

hartnn (hartnn):

can you find the slope of 2x+y = 3 ?

OpenStudy (anonymous):

yes i beloieve so

hartnn (hartnn):

parallel lines have SAME slope, so, the slope of the required line will be ... ?

OpenStudy (anonymous):

i really dont know i think its (5,4)

hartnn (hartnn):

if you believe you can find the slope of 2x+y = 3 , tell me what you got as slope ?

OpenStudy (anonymous):

oh so your basically saying the answer is you cant find a slope with 2x+y=3?

hartnn (hartnn):

no, no , i just want you to find the slope, there is definitely a slope

OpenStudy (anonymous):

The "blueprint" for 2x + y = 3 is also y = mx + b where m = slope.

hartnn (hartnn):

let me give an example, if the equation was 10x+y=44 you can write that as y = -10x +44 comparing with y=mx+c, we get m=slope = -10 so, what will be the slope of 2x+y=3 ?

OpenStudy (anonymous):

i guess y=2x+3 it doesnt even say ah man

OpenStudy (anonymous):

Don't guess, be sure to yourself. If you know the equation is 2x+y=3 and you have to rearrange it for y = mx + b, double check if you did everything right.

hartnn (hartnn):

it would never say, you need to find out, just try to isolate y, its not difficult :) look at my example if you need to

OpenStudy (anonymous):

so basically its y=2x+3?

hartnn (hartnn):

no, when you isolate y, you SUBTRACT 2x from both sides, right ?